# Noncollinear Spin-Flip TDDFT for Potential Energy Surface Crossings: Conical Intersections and Spin Crossings

### v2

@@ Abstract
We recently proposed a scheme to generalize collinear functionals to the noncollinear regime, termed the multicollinear approach. The resulting noncollinear functionals preserve spin symmetry while providing numerically stable higher-order functional derivatives. This scheme has already been applied to noncollinear spin-flip TDDFT and its analytic gradient calculations. In the present work, with the aid of the penalty function method, we employ the noncollinear spin-flip TDDFT in multicollinear scheme to locate potential energy surface crossings. We investigate two distinct types of crossings and analyze their topographical and spin characteristics near the crossing points. The first type is conical intersections, typically involving two singlet states such as the ground and first excited states. The second type involves spin crossings that occur between electronic states with different spin multiplicities, such as between singlet and triplet. These crossing regions enable ultrafast nonadiabatic transitions through either nonadiabatic coupling or spin-orbit coupling, playing a crucial role in photochemistry. Through theoretical analysis and illustrative examples, we demonstrate the advantages of noncollinear spin-flip TDDFT over conventional collinear spin-flip TDDFT or spin-conserving TDDFT. Finally, we systematically evaluate its prospects as an electronic structure method for use in nonadiabatic molecular dynamics.
@@ 1 Introduction
The degeneracy of electronic potential energy surfaces (PESs) is pivotal in determining the dynamics of molecular systems, especially in the context of photochemical reactions and nonradiative transitions1. At the most fundamental level, considering only the intersections between two states, these crossings can be classified into two types according to their spin multiplicities: conical intersections and spin crossings. Conical intersections occur between states of the same spin multiplicity, such as between two singlet states, and serve as photochemical funnels facilitating ultrafast internal conversion2, 3, 4, 5, 6. In contrast, spin crossings, also termed surface intersections, involve states of different spin multiplicities, as exemplified by the intersection between singlet and triplet states, thereby governing intersystem crossing processes via spin-orbit coupling effects7, 8, 9, 10, 11.
The dimensionality of the crossing region is a fundamental aspect in understanding these degeneracies12, 13, 6. This topic traces its origins to the celebrated non-crossing rules of von Neumann and Wigner14, and has been extensively studied in various contexts. For an electronic Hamiltonian representing two diabatic states as a function of nuclear coordinates $\mathbf{R}$:
@@
where $H_{12}=H_{21}$ and the $\times 2$ 2 matrix $H$ is real. The conditions for degeneracy are given by
@@
[…]
Accurate descriptions of conical intersections often require expensive multireference methods, such as multiconfigurational self-consistent field (MCSCF)15 or multireference configuration interaction (MRCI)16. In contrast, computationally efficient methods like density functional theory (DFT)17, 18 and spin-conserving time-dependent DFT (TDDFT)19, 20 fail to produce the correct dimensionality of the intersection between the ground and excited states. The intersections are predicted as $(M-1)$-dimensional, as Eq. 3 is automatically satisfied due to Brillouin’s theorem21. To recover the correct topological features of conical intersections between ground and excited states within DFT and spin-conserving TDDFT, additional modifications are required, such as configuration interaction corrections (CIC)22, 23, dual functional methods (DF)24, particle-particle or hole-hole treatments (pp or hh)25, 26, 27, 28, or the explicit inclusion of double excitations (TDDFT-1D)29.
One promising strategy to restore the correct dimensionality of conical intersections is the adoption of spin-flip methodologies30, 31, 12, 32, 33. According to the spin of the target states, spin-flip TDDFT can be categorized into spin-flip-up and spin-flip-down channels. In this work, we focus on the latter, as shown in Figure 1. The spin-flip TDDFT approach permits a unified treatment of the ground and excited states by introducing single spin-flip excitations from a high-spin reference.
@@
When spin-flip TDDFT was first proposed by Krylov et al.34, it employed collinear exchange-correlation functionals and was therefore called collinear spin-flip TDDFT. The coupling between molecular orbitals of different spins was achieved through the Hartree-Fock (HF) exchange component in hybrid functionals. Collinear spin-flip TDDFT has been widely applied to multireference problems such as diradicals35, 36, triradicals37, 38, photochemical simulations30, 39, 40, 41, and single-molecule magnets42. Despite its practical utility, collinear spin-flip TDDFT has limitations. First, to ensure correct coupling between spin channels, it generally requires the use of hybrid functionals with a substantial fraction of HF exchange, which restricts the applicability of pure functionals. Second, it suffers from the so-called self-splitting problem, where energies of the reference state $|{\mathrm{T}(S_{z}=1)}\rangle$ and the spin-flip state $|{\mathrm{T}(S_{z}=0)}\rangle$ are artificially different, introducing ambiguity and inconsistency. Third, it exhibits spin contamination from three sources: the improper treatment of spin response by collinear functionals, the incompleteness of spin space (Figure 1), and the inherent contamination from the reference triplet state.
@@
On the other hand, noncollinear spin-flip TDDFT, which was initially proposed by Wang and Ziegler43, represents a more rigorous alternative. Within this theoretical framework, pure functionals can directly couple the two spin channels44, making the HF exchange an optional, rather than essential, component. The spin-flip kernel proposed by Wang and Ziegler was initially based on a noncollinear local density approximation (LDA) functional and has been applied to the description of photochemical reactions45. However, when extended to functionals beyond LDA such as generalized gradient approximation (GGA) and meta-GGA, it encounters numerical instabilities due to divisions by zero, which limits its broad applicability44, 46. In our recent work, a general and robust framework for extending noncollinear functionals has been developed, called the multicollinear approach47. This method is applicable to all types of exchange-correlation functionals and provides numerically stable functional derivatives, thereby revitalizing noncollinear spin-flip TDDFT48. Within the multicollinear framework, the calculation of analytic gradients for noncollinear spin-flip TDDFT also becomes simple and natural49. The computation of derivative couplings is currently underway.
@@
In this work, we use noncollinear spin-flip TDDFT in multicollinear approach to locate and characterize conical intersections and spin crossings. Furthermore, we evaluate its potential as an electronic structure computational tool for nonadiabatic dynamics simulations.
@@ 2.1 Spin-Flip TDDFT Formalism
The spin-flip TDDFT is formulated within the linear-response theory. The excitation energies $\Omega$ and corresponding transition amplitudes $X$ are obtained by solving the Casida equation
@@
where $F$ is the Fock matrix, and $K$ is the spin-flip kernel. For simplicity, we will express the $K$ matrix in terms of one-component basis functionals $\{\phi_{\mu}\}$, with the superscripts $\uparrow$ and $\downarrow$ indicating the spin channels. Using GGA-type hybrid functionals as an example, the spin-flip kernel is expressed as
@@
The second term in Eq. 6 originates from the contribution of pure functionals. It naturally appears in noncollinear spin-flip TDDFT but vanishes in collinear spin-flip TDDFT. This term is critically important because it allows the spin-flip kernel to remain nonzero even when pure (nonhybrid) functionals are used. Consequently, the Casida matrix in the left-hand side of Eq. 4 no longer reduces to a diagonal form, enabling the calculation of excitation energies beyond mere orbital energy differences. Within the multicollinear approach, it can be computed as follows
@@
where the collinear second derivatives are averaged with respect to a scaled magnetization $\tilde{m}_{z}=tm_{z}$. As proved by our recent work50, this approach guarantees the correct degeneracy of triplet states and is readily applicable to any functional as long as its second derivatives are available.
@@ 2.2 Locating the Crossing Points
Minimum energy conical intersections (MECIs) between singlet states, as well as minimum energy crossing points (MECPs) for singlet-triplet crossings, correspond to local minima within the seam where two electronic states are degenerate. In this work, we adopt the penalty function scheme originally proposed by Levine et al.,51 in which the following objective function is minimized:
@@
where $I$ and $J$ label the two adiabatic states of interest, $\Delta E_{IJ}=E_{I}-E_{J}$ is their energy gap, $\sigma$ is a penalty prefactor, and $\alpha$ is a smoothing parameter (set to 0.02 a.u. in this work). Optimization of $F_{IJ}$ requires only the energies and energy gradients of the two relevant states, and does not rely on knowledge of their derivative couplings.
Following the protocol in Ref.51, we begin with $\sigma=3.5$ and minimize $F_{IJ}$. If the energy difference $\Delta E_{IJ}$ at the resulting minimum exceeds a preset threshold (0.01 eV in this work), $\sigma$ is increased and the minimization is repeated iteratively until the degeneracy condition is satisfied. The optimization follows the convergence criteria for both step size and gradient norm of the objective function, as stated in Ref.51.
@@ 3 Noncollinear Spin-Flip TDDFT for PES Crossings
In this section, we present four examples illustrating the ability of noncollinear spin-flip TDDFT to locate PES crossings using the PySCF package52, 53. In the first application, $\mathrm{S}_{0}/\mathrm{S}_{1}$ conical intersections are located, which is arguably the most important class of crossings. Conventional DFT combined with spin-conserving TDDFT often fails to produce the correct topological features in such cases. Noncollinear spin-flip TDDFT not only successfully identifies these intersections with the correct topological structure, but also significantly reduces the spin contamination present in collinear spin-flip TDDFT. In the second application, we demonstrate that collinear spin-flip TDDFT yields artificial PES crossings when pure functionals are used. In contrast, noncollinear spin-flip TDDFT produces reasonable results and avoids such spurious features. The third and fourth applications focus on locating $\mathrm{S}_{0}/\mathrm{T}$ and $\mathrm{S}_{1}/\mathrm{T}$ crossings, respectively. Thanks to the zero excitation energy theorem50, noncollinear spin-flip TDDFT yields consistent triplet energies, enabling an unambiguous identification of singlet-triplet intersections.
@@ 3.1 $\mathrm{S}_{0}/\mathrm{S}_{1}$ Conical Intersections
Due to the inability of DFT and spin-conserving TDDFT to correctly describe the topological features of the $\mathrm{S}_{0}/\mathrm{S}_{1}$ conical intersection, researchers have often turned to collinear spin-flip TDDFT to address this problem. Previous studies have shown that collinear spin-flip TDDFT typically locates a reasonable geometry for the intersection30, 31, 12, 32, 33. However, the computed $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ states sometimes exhibit a certain degree of spin contamination. Spin contamination leads to ambiguities in the spin multiplicity of the conical intersection, thereby complicating its interpretation and analysis. It can also adversely affect subsequent calculations, such as derivative couplings.
To investigate how noncollinear spin-flip TDDFT handles this issue, we revisit the computational examples from Ref. 54. We select three molecular systems, including ethylene, penta-2,4-dieniminium cation (PSB3), and styrene, from Ref. 54, in which collinear spin-flip TDDFT exhibited significant spin contamination. MECI geometries are optimized for these three molecules at the BHHLYP/6-31+G**55, 56, 57, 58, 59 level, as shown in Figure 2a. The resulting structures are similar to those obtained from collinear spin-flip TDDFT and the high-level MRCISD calculations. Detailed data are provided in Table 1.
@@
Figure 2: (a) Geometries of MECIs obtained by MRCISD (gray, from Ref. 54), collinear spin-flip TDDFT (green, from Ref. 54), and noncollinear spin-flip TDDFT (red, this work). Structures are visualized using PyMOL60. (b) PESs of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ near the MECI of ethylene. The twisting angles follow the convention in Ref. 30.
@@
Table 1: Root Mean Square Deviations (RMSD, in $\mathrm{\SIUnitSymbolAngstrom}$) of MECI geometries with respect to MRCISD results54 and $\langle{\hat{S}^{2}}\rangle$ values of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ states at the MECI geometries.
Collinear BHHLYPa Noncollinear BHHLYP Noncollinear PBE
System RMSD $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{0}}$ $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{1}}$ RMSD $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{0}}$ $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{1}}$ RMSD $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{0}}$ $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{1}}$
Ethylene 0.077 0.277 0.435 0.019 0.069 0.078 0.191 0.011 0.029
PSB3 0.240 0.092 0.579 0.048 0.071 0.107 0.335 0.033 0.027
Styrene 0.102 0.072 0.932 0.101 0.120 0.121 0.116 0.022 0.029
aTaken from Ref. 54.
@@
To verify that noncollinear spin-flip TDDFT indeed produces the correct topological features, we take the ethylene molecule as an example and plot the $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ PESs near the conical intersection point using B3LYP/6-31+G**55, 61. As shown in Figure 2b, the surfaces indeed exhibit the characteristic conical intersection topology.
To examine the spin contamination in the $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ states computed by noncollinear spin-flip TDDFT, we report the $\langle{\hat{S}^{2}}\rangle$ values at the MECIs in Table 1. The results show that when both collinear and noncollinear spin-flip TDDFT employ the BHHLYP functional, the latter exhibits a reduction in spin contamination by approximately half an order of magnitude. Moreover, noncollinear spin-flip TDDFT not only works well with hybrid functionals but also yields accurate structures and low spin contamination when used with pure functionals such as PBE62, demonstrating its robustness across different types of functionals.

### v3

# Contributed $\altaffiliationContributed equally to this work$

@@ Abstract
We recently proposed a scheme to generalize collinear functionals to the noncollinear regime, termed the multicollinear approach, which preserves spin symmetry and provides numerically stable higher-order functional derivatives. This scheme has been applied to noncollinear spin-flip TDDFT and its analytic gradient calculations. In the present work, with the aid of the penalty function method, we employ the noncollinear spin-flip TDDFT in multicollinear scheme to locate potential energy surface crossings. We investigate two distinct types of crossings and analyze their topographical and spin characteristics near the crossing points. The first type is conical intersections between states of the same spin multiplicity, such as the singlet ground and excited states. The second type involves spin crossings that occur between electronic states with different spin multiplicities, such as between singlet and triplet. These crossing regions enable ultrafast nonadiabatic transitions through either nonadiabatic coupling or spin-orbit coupling, playing a crucial role in photochemistry. Through theoretical analysis and illustrative examples, we demonstrate the advantages of noncollinear spin-flip TDDFT over conventional collinear spin-flip TDDFT or spin-conserving TDDFT. Finally, we systematically evaluate its prospects as an electronic structure method for use in nonadiabatic molecular dynamics.
@@ 1 Introduction
The degeneracy of electronic potential energy surfaces (PESs) is pivotal in determining the dynamics of molecular systems, especially in the context of photochemical reactions and nonradiative transitionsschuurman2018dynamics. At the most fundamental level, considering only the intersections between two states, these crossings can be classified into two types according to their spin multiplicities: conical intersections and spin crossings. Conical intersections occur between states of the same spin multiplicity, such as between two singlet states, and serve as photochemical funnels facilitating ultrafast internal conversionyarkony1996diabolical, yarkony1999s1s0, domcke2012role, haas2014conical, matsika2021electronic. In contrast, spin crossings involve states of different spin multiplicities, as exemplified by the intersection between singlet and triplet states, thereby governing intersystem crossing processes via spin-orbit coupling effectsharvey1998singlet, casavecchia2015reaction, vanuzzo2016reaction, li2019intersystem, wang2021conical.
The dimensionality of the crossing region is a fundamental aspect in understanding these degeneraciesgozem2014shape, ryabinkin2017geometric, matsika2021electronic. This topic traces its origins to the celebrated non-crossing rules of von Neumann and Wignernewmann1929behaviour, and has been extensively studied in various contexts. As discussed by Meadmead1979noncrossing, the dimensionality of the intersection depends on whether spin-orbit coupling is present and on whether the system has an even or odd number of electrons. For the simplest case of a non-relativistic, even-electron system, the electronic Hamiltonian representing two diabatic states as a function of nuclear coordinates $\mathbf{R}$ is given by:
@@
where $H_{12}=H_{21}$ and the $2\times 2$ matrix $H$ is real. The conditions for degeneracy are given by
@@
which define a subspace where two PESs are degenerate. For a molecule with $N$ atoms, the number of internal nuclear degrees of freedom is M=3N6M=3N-6 (or 3N53N-5 for linear molecules). To simultaneously satisfy the two independent conditions above, two degrees of freedom are constrained, and thus the intersection forms an (M2)(M-2)-dimensional seam, known as a conical intersection. The remaining two dimensions, referred to as the branching space, lift the degeneracy. In contrast, spin crossings are described by an (M1)(M-1)-dimensional intersection, as degeneracy is lifted in only one coordinate direction, because Eq. 3 is always satisfied for states of different spin multiplicities.
Accurate descriptions of conical intersections often require expensive multireference methods, such as multiconfigurational self-consistent field (MCSCF)helgaker2000 or multireference configuration interaction (MRCI)shavitt1977. In contrast, computationally efficient methods like density functional theory (DFT)hohenberg1964inhomogeneous, kohn1965selfconsistent and spin-conserving time-dependent DFT (TDDFT)runge1984densityfunctional, casida1995timedependent fail to produce the correct dimensionality of the intersection between the ground state (using DFT) and excited state (using TDDFT), as Eq. 3 is automatically satisfied due to Brillouin’s theoremlevine2006conical. To recover the correct topological features of conical intersections between ground and excited states within DFT and spin-conserving TDDFT, additional modifications are required, such as configuration interaction corrections (CIC)li2014configuration, xu2025conical, dual functional (DF) methodsshu2017dualfunctional, particle-particle (pp) or hole-hole (hh) treatmentsyang2013double, vanaggelen2013exchangecorrelation, yang2016conical, bannwarth2020hole, or the explicit inclusion of double excitations (TDDFT-1D)teh2019simplest.
One promising strategy to restore the correct dimensionality of conical intersections is the adoption of spin-flip methodologiesminezawa2009optimizing, harabuchi2013automated, gozem2014shape, zhang2014analytic, winslow2020comparison. Depending on the type of excitation, spin-flip TDDFT can be categorized into spin-flip-up and spin-flip-down channels. In this work, we focus on the latter, as shown in Figure 1. By introducing single spin-flip-down excitations from a high-spin reference state $|{\mathrm{T}(S_{z}=1)}\rangle$, the spin-flip-down TDDFT approach enables a unified description of both the ground and excited states.
@@
When spin-flip TDDFT was first proposed by Krylov et al.shao2003spin, it employed collinear exchange-correlation functionals and was therefore called collinear spin-flip TDDFT. In such framework, the coupling between molecular orbitals of different spins arises solely from the Hartree-Fock (HF) exchange component in hybrid functionals. Collinear spin-flip TDDFT has since been widely applied to diverse systems, such as diradicalskishi2007finitefield, hossain2017photoelectron, triradicalscristian2004bonding, krylov2005triradicals, photochemical simulationsminezawa2009optimizing, harabuchi2014dynamics, yue2018performance, zhang2021nonadiabatic, and single-molecule magnetsvalero2011magnetic. Despite its practical utility, collinear spin-flip TDDFT has several limitations. First, to ensure proper coupling between spin channels, it generally requires hybrid functionals with a large fraction of HF exchange, which restricts the applicability of pure functionals. Second, it suffers from the so-called self-splitting problem, where energies of the reference state $|{\mathrm{T}(S_{z}=1)}\rangle$ and the spin-flip state $|{\mathrm{T}(S_{z}=0)}\rangle$ differ artificially, introducing ambiguity and inconsistency in determining the triplet energies. Third, it is prone to spin contamination from three sources: the incomplete spin treatment inherent in the collinear spin-flip kernel, the incompleteness of spin space (Figure 1), and the spin contamination already present in the high-spin reference state.
@@
On the other hand, noncollinear spin-flip TDDFT has a clearer and more rigorous theoretical foundation. Wang and Zieglerwang2004timedependent realized that applying the nonrelativistic version of two-component noncollinear TDDFTgao2005timedependent to a collinear system naturally leads to two decoupled equations: one is the conventional spin-conserving TDDFT equation (which only requires a collinear functional), and the other is the noncollinear spin-flip TDDFT equation (which requires a noncollinear functional). Ref. li2025analytic provides a detailed derivation of this decoupling. A key implication of this framework is that pure functionals can directly couple the two spin channelsbernard2012general, making the HF exchange an optional, rather than mandatory, component. The spin-flip kernel originally proposed by Wang and Ziegler was based on a noncollinear local density approximation (LDA) functional and has been applied to the description of photochemical reactionshuix-rotllant2010assessment. However, when extended to functionals beyond LDA such as generalized gradient approximation (GGA) and meta-GGA, it encounters numerical instabilities due to divisions by zero, which limits its broad applicabilitybernard2012general, li2012theoretical. In our recent work, a general and robust framework for extending noncollinear functionals has been developed, called the multicollinear approachpu2023noncollinear. This method is applicable to all types of exchange-correlation functionals and provides numerically stable functional derivatives, thereby revitalizing noncollinear spin-flip TDDFTli2023noncollinear. Within the multicollinear framework, the calculation of analytic gradients for noncollinear spin-flip TDDFT also becomes simple and naturalli2025analytic. The computation of derivative couplings is currently underway.
@@
In this work, noncollinear spin-flip TDDFT within the multicollinear approach is used to locate and characterize conical intersections and spin crossings. Furthermore, we evaluate its potential as an electronic structure computational tool for nonadiabatic dynamics simulations.
@@ 2.1 Spin-Flip TDDFT Formalism
In this section, we present the working equations for noncollinear spin-flip TDDFT. Detailed derivations can be found in Refs. li2025analytic, wang2025zero.
In the spin-flip-down formalism, the excitation energies $\Omega$ and the corresponding transition amplitudes $X$ are obtained by solving the Casida equation:
@@
where $F$ is the Fock matrix, and $K$ is the spin-flip kernel. For simplicity, we will express the $K$ matrix in terms of one-component basis functionals $\{\phi_{\mu}\}$, with the superscripts $\uparrow$ and $\downarrow$ indicating the spin channels. The spin-flip kernel is expressed as
@@
The second term in Eq. 6 originates from the contribution of pure functionals. It naturally appears in noncollinear spin-flip TDDFT but vanishes in collinear spin-flip TDDFT. This term is critically important because it allows the spin-flip kernel to remain nonzero even when pure functionals are used. Consequently, the Casida matrix in the left-hand side of Eq. 4 no longer reduces to a diagonal form, enabling the calculation of excitation energies beyond mere orbital energy differences. Within the multicollinear approach, it can be computed as followsli2023noncollinear
@@
where the collinear second derivatives are averaged with respect to a scaled magnetization $\tilde{m}_{z}=tm_{z}$. As proved by our recent workwang2025zero, this approach guarantees the correct degeneracy of triplet states and is readily applicable to any collinear functional as long as its second derivatives are available.
@@ 2.2 Locating the Crossing Points
Minimum energy conical intersections (MECIs) between singlet states, as well as minimum energy crossing points (MECPs) for singlet-triplet crossings, correspond to local minima within the seam where two electronic states are degenerate. In this work, we adopt the penalty function scheme originally proposed by Levine et al.,levine2008optimizing in which the following objective function is minimized:
@@
where $I$ and $J$ label the two adiabatic states of interest, $\Delta E_{IJ}=E_{I}-E_{J}$ is their energy gap, $\sigma$ is a penalty prefactor, and $\alpha$ is a smoothing parameter (set to 0.02 a.u. in this work). Optimization of $F_{IJ}$ requires only the energies and energy gradients of the two relevant states, without the need for derivative couplings.
Following the protocol in Ref.levine2008optimizing, we begin with $\sigma=3.5$ and minimize $F_{IJ}$. If the energy difference $\Delta E_{IJ}$ at the resulting minimum exceeds a preset threshold (0.01 eV in this work), $\sigma$ is increased and the minimization is repeated iteratively until the degeneracy condition is satisfied. The optimization follows the convergence criteria for both step size and gradient norm of the objective function, as stated in Ref.levine2008optimizing.
@@ 3 Noncollinear Spin-Flip TDDFT for PES Crossings
In this section, we illustrate the capabilities of noncollinear spin-flip TDDFT for investigating PES crossings through several representative applications. First, we locate $\mathrm{S}_{0}/\mathrm{S}_{1}$ conical intersections, which is arguably the most important class of crossings. Noncollinear spin-flip TDDFT successfully identifies these intersections with the correct topological structure and significantly reduces the spin contamination observed in collinear spin-flip TDDFT. Second, we show that collinear spin-flip TDDFT can produce artificial PES crossings when pure functionals are used, whereas the noncollinear approach avoids such spurious features. Third and fourth, we consider $\mathrm{S}_{0}/\mathrm{T}$ and $\mathrm{S}_{1}/\mathrm{T}$ crossings. With the zero excitation energy theoremwang2025zero, noncollinear spin-flip TDDFT provides consistent triplet energies, enabling unambiguous identification of singlet-triplet intersections. Finally, we demonstrate its versatility by describing $\mathrm{D}_{0}/\mathrm{D}_{1}$ conical intersection in an odd-electron system. All calculations are performed using the PySCF packagesun2020recent, pyscfforge.
@@ 3.1 $\mathrm{S}_{0}/\mathrm{S}_{1}$ Conical Intersections
Collinear spin-flip TDDFT has often been employed to locate $\mathrm{S}_{0}/\mathrm{S}_{1}$ conical intersections, and previous studies have shown that it generally provides reasonable MECI geometriesminezawa2009optimizing, harabuchi2013automated, gozem2014shape, zhang2014analytic, winslow2020comparison. However, the resulting $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ states can exhibit significant spin contamination. This contamination introduces ambiguities in the spin multiplicity of the intersection and can adversely affect subsequent calculations, such as derivative couplings.
Here, we revisit the benchmark systems of Ref. nikiforov2014assessment. Three representative molecules–ethylene, penta-2,4-dieniminium cation (PSB3), and styrene–are selected, all of which were previously shown to suffer from significant spin contamination when treated with collinear spin-flip TDDFT. Here, MECIs are optimized for these systems using the BHHLYP functionalbecke1988densityfunctional, lee1988development, becke1993densityfunctional with the 6-31+G** basis setditchfield1971selfconsistent, hehre1972self, hariharan1973influence, clark1983efficient, following the same level of theory as in Ref. nikiforov2014assessment to ensure direct comparability. The resulting MECI geometries from noncollinear spin-flip TDDFT, collinear spin-flip TDDFT, and MRCISD calculations are illustrated in Figure 2a, with the detailed results summarized in Table 1.
With the same BHHLYP functional, noncollinear spin-flip TDDFT yields structures that are in closer agreement with high-level MRCISD results than those from the collinear approach. It is worth noting that the $\langle{\hat{S}^{2}}\rangle$ values are substantially reduced by approximately half an order of magnitude for both $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ states. These results highlight the effectiveness of noncollinear spin-flip TDDFT in accurately locating conical intersections while mitigating spin contamination issues.
@@
Figure 2: (a) Geometries of MECIs obtained by MRCISD (gray, from Ref. nikiforov2014assessment), collinear spin-flip TDDFT with BHHLYP (green, from Ref. nikiforov2014assessment), and noncollinear spin-flip TDDFT with BHHLYP (red, this work). Structures are visualized using PyMOLPyMOL. (b) PESs of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ near the MECI of ethylene. The twisting angles follow the convention in Ref. minezawa2009optimizing. Structures are visualized using VMDhumphrey1996vmd, stone1998tachyon.
@@
Table 1: Root Mean Square Deviations (RMSDs, in $\mathrm{\SIUnitSymbolAngstrom}$) of MECI geometries with respect to MRCISD resultsnikiforov2014assessment and $\langle{\hat{S}^{2}}\rangle$ values of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ states at the MECI geometries.
Collinear BHHLYPa Noncollinear BHHLYP
System RMSD $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{0}}$ $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{1}}$ RMSD $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{0}}$ $\langle{\hat{S}^{2}}\rangle_{\mathrm{S}_{1}}$
Ethylene 0.077 0.277 0.435 0.018 0.023 0.092
PSB3 0.240 0.092 0.579 0.035 0.071 0.107
Styrene 0.102 0.072 0.932 0.031 0.113 0.117
aTaken from Ref. nikiforov2014assessment.
Importantly, the noncollinear spin-flip TDDFT framework is not limited to hybrid functionals. It can be combined with various types of exchange-correlation functionals, including LDA, GGA, meta-GGA, and hybrid functionals. Appendix A presents detailed tests on the performance of these different functionals, as well as a discussion of the basis set dependence of the MECI geometries.
@@
To verify that noncollinear spin-flip TDDFT indeed captures the topological features of the intersection, we take the ethylene molecule as an example and plot the $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ PESs near the conical intersection point. As shown in Figure 2b, the computed potential energy surfaces clearly display the characteristic double-cone topology, demonstrating that the noncollinear approach reproduces the correct intersection shape. The B3LYP functional is employed here as a representative example, though other functionals yield similar conical intersection structures.